OpenStudy (jenniferjuice):

Given that f(x) = 2x + 1 and g(x) = −5x + 2, solve for f(g(x)) when x = 3. −91 −33 −25 35

OpenStudy (anonymous):

Okay so we'll solve g(x) first. g(3) = -5(3) + 2 [Substitute]

OpenStudy (anonymous):

f(3)=2x+1 g(3)=-5x+2

OpenStudy (anonymous):

g(3) = -15 + 2 g(3) = -13

OpenStudy (muzzack):

yes @jcpd910 so solve the equation now you have -13

OpenStudy (anonymous):

g(3) will replace the x in f(x)

OpenStudy (muzzack):

yes replace it easy i think @jenniferjuice gets it

OpenStudy (anonymous):

Let's see if she can solve it now.

OpenStudy (zzr0ck3r):

\(f(g(x)) = 2(-5x+2)+1\)

OpenStudy (anonymous):

zzr0ck3r we are not giving out answers, we are helping people make their own.

OpenStudy (zzr0ck3r):

lol tnx man, the 99 shuold tell you I have done this before....

OpenStudy (zzr0ck3r):

I gave no answers. I showed how to compose the function...

OpenStudy (anonymous):

I think you made it too easy. A lot of people don't understand how g(x) replaces the x in f(x) when you're solving for f(g(x))

OpenStudy (anonymous):

But let's not bicker over petty problems. :D

OpenStudy (zzr0ck3r):

well there are two ways of thinking about this. One: find the answer to g(a) and pluf that into f(x) Two: compose the function f(g(x)) then evaluate f(g(a)). I like to make sure people see both options

OpenStudy (anonymous):

I did not get 17. Can I see your work please?

OpenStudy (jenniferjuice):

f(g(x))=2(−5x+2)+1 =2(−5*3+2)+1 = 2(-2+2)+1 2(0+2)+1 0+2+1 3

OpenStudy (anonymous):

-5*3 is not -2.

OpenStudy (jenniferjuice):

-25?

OpenStudy (anonymous):

2(-5*3+2)+1 2(-15+2)+1

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

* means multiplied by.

OpenStudy (jenniferjuice):

-25?

OpenStudy (anonymous):

You got it. :)